# EAGLE: The First Event Camera Dataset Gathered by an Agile Quadruped Robot

### v1

@@ Abstract
When legged robots perform agile movements, traditional RGB cameras often produce blurred images, posing a challenge for accurate state estimation. Event cameras, inspired by biological vision mechanisms, have emerged as a promising solution for capturing high-speed movements and coping with challenging lighting conditions, owing to their significant advantages, such as low latency, high temporal resolution, and a high dynamic range. However, the integration of event cameras into agile-legged robots is still largely unexplored. Notably, no event camera-based dataset has yet been specifically developed for dynamic legged robots. To bridge this gap, we introduce EAGLE (Event dataset of an AGile LEgged robot), a new dataset comprising data from an event camera, an RGB-D camera, an IMU, a LiDAR, and joint angle encoders, all mounted on a quadruped robotic platform. This dataset features more than 100 sequences from real-world environments, encompassing various indoor and outdoor environments, different lighting conditions, a range of robot gaits (e.g., trotting, bounding, pronking), as well as acrobatic movements such as backflipping. To our knowledge, this is the first event camera dataset to include multi-sensory data collected by an agile quadruped robot.
@@
Event camera dataset, Agile quadruped robot, Multi-sensory data.
@@ I INTRODUCTION
In this paper, we introduce a comprehensive dataset tailored for quadruped robots that can perform dynamic movements in various lighting conditions and environments. The superior mobility of quadruped robots holds significant potential in applications requiring rapid navigation in complex environments including confined spaces and rough terrains. However, despite significant advancement in their locomotion capability [1, 2], their application remains limited to the scenarios without immediate needs or rapid exploration [3, 4]. This constraint delays the broader adoption of the extremely capable legged robotic system to perform hazardous and time-sensitive tasks, such as disaster response, search and rescue, and firefighting [5, 6]. One of the most pressing issues is the compromised perception capability during swift movements: motion blur in RGB images and distortion in LiDAR point clouds hinder precise pose estimation and environmental understanding.
@@
Depth Sensor | IMU | Platform | Environment | Ground-truth Pose | ||
| D-eDVS [7] | RGB-D | ✗ | Handheld | Indoor | MoCap | |||
| evbench [8] | RGB-D | ✗ | Wheeled robot | Indoor | Odometer | |||
| Mueggler et al. [9] | ✗ | ✓ | Handheld | Indoor/Outdoor | MoCap | |||
| MVSEC [10] | LiDAR-16 | ✓ | Multiple Robots | Indoor/Outdoor | MoCap/Cartographer | |||
| UZH-FPV [11] | ✗ | ✓ | Drone | Indoor/Outdoor | MoCap | |||
| ViViD [12] | RGB-D/LiDAR-16 | ✓ | Handheld | Indoor/Outdoor | MoCap/LeGO-LOAM | |||
@@
RGB-D/LiDAR-64 | ✓ | Handheld/Car | Indoor/Outdoor | MoCap/LeGO-LOAM/RTK-GPS | ||
| DSEC [14] | LiDAR-16 | ✗ | Car | Outdoor | RTK-GPS | |||
| AGRI-EBV [15] | RGB-D/LiDAR-16 | ✓ | Wheeled Robot | Outdoor | LeGO-LOAM | |||
| TUM-VIE [16] | ✗ | ✓ | Handheld/Bike | Indoor/Outdoor | MoCap | |||
| VECtor [17] | RGB-D/LiDAR-128 | ✓ | Handheld/Scooter | Indoor | MoCap/ICP | |||
| M3ED [18] | LiDAR-64 | ✓ | Multiple Robots | Indoor/Outdoor | Faster-LIO | |||
| EAGLE |
@@
* Extension -16 represents a 16-beam LiDAR.
@@
Event cameras, similar to the human eye, detect logarithmic intensity changes in images and offer low latency and high temporal resolution, making them exceptionally suited for handling rapid robotic movements. Additionally, their high dynamic range (HDR) effectively addresses challenges in poorly-lit and brightly-lit settings, thereby broadening potential applications for robots [19, 20]. Recent studies on event cameras have increasingly gained attention in systems requiring rapid perception, such as drones and autonomous vehicles [21, 22], yet their use has not been widely recognized within the legged robot community. Once properly configured, these unique sensors can significantly enhance state estimation and terrain perception, greatly expanding the capabilities of legged robots.
@@
While promising, event cameras also have drawbacks when used as standalone sensing systems. Due to their operating mechanism of detecting changes in brightness, event data inherently depend on the motion of the camera itself (ego-motion), making it challenging to differentiate between the camera’s movement and changes in the external environment. This issue can obscure the clarity of visual data representations, especially when the robot moves slowly or in directions parallel to environmental textures. Overcoming these limitations and enhancing the functionality of agile legged robots in diverse environments necessitates the development of algorithms for the synergistic integration of multimodal sensors. However, as of 2024, there is a lack of comprehensive datasets to support this research.
@@
To bridge this gap, we introduce a new dataset collected by a unique sensor suite on an agile quadruped robot, MIT Mini-Cheetah [23]. Our system features an event camera, an RGB-D camera, an IMU, a LiDAR, and joint angle sensors. The integration of event and RGB cameras provides the legged robot with enhanced capabilities for both slow and rapid motions. Depth information, crucial for accurate state estimation and navigation, is supplied by the RGB-D camera and LiDAR. The IMU, critical for capturing high-frequency angular and acceleration data, is inherently immune to disturbances from external dynamic objects. Furthermore, joint angle data can also provide odometry information through forward kinematics.
@@
Our dataset includes various distinct robot gaits, including trotting, bounding, and pronking. It also encompasses a wide range of daily indoor and outdoor environments, under varying lighting conditions (e.g., daytime, nighttime, well-lit, dark, high dynamic range, and blinking light conditions). This diverse collection provides an in-depth perspective on the robot’s capabilities under varying conditions and times. Throughout the dataset building, we aim to contribute to the enhancement of the legged robots’ resilience and adaptability across a multitude of real-world environments.
In summary, the main contributions of this paper include:
We present the first dataset that provides a comprehensive collection of sensor data from an agile quadruped robot, including inputs from an event camera, RGB-D camera, IMU, LiDAR, and joint encoders.
We provide data across 31 distinct environments, within each of which we captured data under varying lighting conditions and through diverse robotic locomotion gaits.
We offer 6 DoF ground-truth poses from a motion capture system or an advanced SLAM algorithm, complemented with precise intrinsic, extrinsic, and temporal synchronization parameters.
All the data sequences are publicly available at https://daroslab.github.io/EAGLE/.
@@ II RELATED WORK
The growing interest in event cameras has led to numerous datasets aimed at various research goals. In this section, we provide an overview of these event camera datasets, along with a comparative summary in Table I.
@@
The first event camera dataset, introduced by Weikersdorfer et al. in 2014 [7], featured a eDVS event camera [24] and an RGB-D camera. In 2016, Barranco et al. [8] utilized an event camera () for visual navigation of a mobile wheeled robot. Mueggler et al. [9], in 2017, released a comprehensive handheld dataset that includes motion speed information as a reference for investigating the capabilities of an event camera (). These datasets have played a crucial role in the introduction of event cameras in research communities, yet the resolution of the event cameras used in these early stages was relatively low.
@@
Subsequent developments have led to datasets with higher-resolution event cameras. In 2022, Lee et al. expanded the ViViD dataset [12] into ViViD++ [13], incorporating a high-resolution event camera () for driving scenarios. This update included repeated data collection along the predefined trjectories at different times, categorizing motions as slow, unstable, or aggressive. Klenk et al. [16] introduced the TUM-VIE dataset with stereo event camera recordings (), capturing diverse environments and activities such as walking, running, skating, and biking. Additionally, Gao et al. [17] released the VECtor dataset, which includes head-mounted and pole-mounted stereo event camera sequences () in indoor settings. These enhanced-resolution event camera datasets improved image clarity and details, enabling more precise analysis.
@@
In the evolution of event camera datasets, datasets for specific research purposes have also been developed. The DDD17 [25] and DDD20 [26] datasets, for instance, provide comprehensive metadata including vehicle speed, GPS positioning, and detailed driving dynamics such as steering, throttle, and brake inputs, improving steering prediction accuracy. Fischer et al. [27] aimed to tackle the place recognition challenge by mounting a DAVIS346 event camera () on a car, capturing data along the same trajectory at different times. Zujevs et al. [15] created agricultural robotics datasets in varied agricultural settings using a mobile wheeled robot. Additionally, Gehrig et al. [14] utilized a dual stereo camera setup with a 16-channel Velodyne LiDAR and an RTK-GPS system for automotive driving scenarios. These datasets delve deeper into specific fields and provide the opportunity to explore the potential of event cameras in focused applications.
@@
Another trend in the development of event camera datasets is their application in various robotic platforms. Delmerico et al. [11] introduced a dataset focused on aggressive drone racing motions. Zhu et al. [10] launched the first multi-robot dataset in 2018, featuring indoor and outdoor places with handheld devices, hexacopters, vehicles, and motorcycles. Additionally, Chaney et al. [18] unveiled the M3ED dataset, showcasing event cameras across diverse sensor arrays on multiple robotic platforms, including unmanned aerial vehicles, wheeled ground vehicles, and legged robots. These datasets further explore the unique capabilities of event cameras across different robotic platforms.
Although the M3ED dataset explores event cameras with a quadruped robot Spot, it does not fully investigate the dynamic capabilities unique to quadruped robots, such as various gaits like trotting, pronking, and bounding. The M3ED dataset utilizes the trotting gait only within relatively slow and conservative operating regions without involving dynamic movements. In addition, the M3ED dataset only contains sparse depth images generated from point cloud maps. On the other hand, our dataset includes dense depth data, various quadruped robot gait motions, and navigation in various environments under different lighting conditions.
@@ III-A Sensors Overview
The DAVIS346 event camera ( resolution), with a horizontal and vertical field of view and a dB dynamic range, includes a built-in 6-axis IMU and the capability to capture RGB images. The DVXplorer Lite ( resolution), featuring a horizontal and vertical field of view and a dB dynamic range, also includes a built-in 6-axis IMU but cannot output RGB images. Both cameras feature sub-millisecond latency. We put an infrared filter in front of both event cameras to block emissions from the RealSense RGB-D camera.
The RealSense camera D455 operates at resolution and , with and fields of view for its RGB and depth cameras, respectively. Timestamps are set at the midpoint of exposure, and auto exposure is enabled to obtain well-exposed images in different lighting conditions. Thus, the timestamp gap between RGB and depth frames varies, typically from to . VectorNav VN-100 is a 9-axis IMU that comprises a 3-axis gyroscope, 3-axis accelerometer, and 3-axis magnetometer at . Additionally, the dataset also includes the orientation data produced by the IMU sensor in quaternion representation. The Velodyne VLP-16 LiDAR delivers highly accurate depth measurements as point clouds at . The LiDAR is mounted atop the Mini Cheetah, ensuring an unobstructed view of the surrounding environment. The Mini Cheetah’s 12 custom motors feature joint encoders that measure angular position at . For this dataset, data is recorded at .
@@ III-C Time Synchronization
Every time we start the data collection, the Mini-cheetah performs a body pitching swing to produce specific sinusoidal patterns in the sensor data. Since each camera sensor includes built-in IMU sensors, we compare the angular velocity data from the IMUs in the event and RealSense cameras with that of the VectorNav IMU. During this process, an event camera is used as a reference, and the timestamps of other sensors are offset based on it. Then, the robot’s joint encoder data are synchronized with the event cameras by comparing the angular velocities of the knee joint and the built-in IMU of the event camera.
@@ IV Dataset
The EAGLE dataset contains 106 sequences including 50 indoor, 40 outdoor, and 16 backflip sequences. Our dataset aims to evaluate algorithm performance: 1) across diverse quadruped gaits and lighting conditions via indoor and outdoor sequences, and 2) in highly dynamic motions through backflip sequences. The name of each sequence reflects the recording environment, lighting conditions, and the gaits of the quadruped robot. For instance, downtown1_day_trot represents data gathered in a downtown environment during daytime with a trot gait, while classroom_blinking_comb indicates data collected in a classroom environment under blinking light conditions with a combination of trotting, bounding, and pronking gaits. Fig. 4 further illustrates dynamic characteristics inherent to each gait. In addition, we marked the initial positions of the robot’s four feet in each sequence and confirmed its return to these exact positions, providing an additional metric for pose estimation.
@@ IV-A Outdoor Sequences
The outdoor sequences were captured in 10 distinctive outdoor environments, with four sequences per environment. Each set of four sequences includes variations of two different sets of quadruped gaits (trot-only and combined gaits) under different periods of the day (daytime and nighttime), ensuring a comprehensive representation of real-world settings. Fig. 5 provides an illustration of all outdoor environments and associated challenges such as varied light conditions from overexposure to underexposure, sunny, cloudy, and foggy weather, diverse visual patterns including featureless terrains and distant skies, dynamic objects (e.g., vehicles and pedestrians), and slippery surfaces. The inclusion of these elements in our dataset mirrors a wide range of real-world challenges, broadening the evaluation spectrum and offering a comprehensive assessment opportunity.

### v3

# CEAR: Comprehensive Event Camera Dataset for Rapid Perception of Agile Quadruped Robots

@@ Abstract
When legged robots perform agile movements, traditional RGB cameras often produce blurred images, posing a challenge for rapid perception. Event cameras have emerged as a promising solution for capturing rapid perception and coping with challenging lighting conditions thanks to their low latency, high temporal resolution, and high dynamic range. However, integrating event cameras into agile-legged robots is still largely unexplored. To bridge this gap, we introduce CEAR, a dataset comprising data from an event camera, an RGB-D camera, an IMU, a LiDAR, and joint encoders, all mounted on a dynamic quadruped, Mini Cheetah robot. This comprehensive dataset features more than 100 sequences from real-world environments, encompassing various indoor and outdoor environments, different lighting conditions, a range of robot gaits (e.g., trotting, bounding, pronking), as well as acrobatic movements like backflip. To our knowledge, this is the first event camera dataset capturing the dynamic and diverse quadruped robot motions under various setups, developed to advance research in rapid perception for quadruped robots. The CEAR dataset is available at https://daroslab.github.io/cear/.
@@
Data Sets for SLAM, Data Sets for Event Camera, Data Sets for Robot Learning, Sensor Fusion.![[Uncaptioned image]](x1.png)
@@ I INTRODUCTION
In this paper, we introduce a comprehensive dataset developed for rapid perception of agile quadruped robots that can perform dynamic movements in various lighting conditions and environments. The superior mobility of quadruped robots holds significant potential in applications requiring rapid navigation in complex environments. However, despite significant advancement in their locomotion capability [1, 2], their application remains limited to the scenarios without immediate needs or rapid exploration [3, 4]. This constraint delays the broader adoption of the extremely capable legged robotic system to perform hazardous and time-sensitive tasks, such as disaster response, search and rescue, and firefighting [5, 6]. One of the most pressing issues is the compromised perception capability during swift movements: motion blur in RGB images and distortion in LiDAR point clouds hinder precise pose estimation and environmental understanding.
@@
Depth Sensor | IMU | Platform | Environment | Ground-truth Pose | ||||
D-eDVS [7] | 128px 128px | RGB-D | ✗ | Handheld | Indoor | MoCap | ||||
evbench [8] | 240px 180px | RGB-D | ✗ | Wheeled robot | Indoor | Odometer | ||||
Mueggler et al. [9] | 240px 180px | ✗ | ✓ | Handheld | Indoor/Outdoor | MoCap | ||||
MVSEC [10] | 2 346px 260px | LiDAR-16 | ✓ | Multiple Robots | Indoor/Outdoor | MoCap/Cartographer | ||||
UZH-FPV [11] | 346px 260px | ✗ | ✓ | Drone | Indoor/Outdoor | MoCap | ||||
ViViD [12] | 240px 180px | RGB-D/LiDAR-16 | ✓ | Handheld | Indoor/Outdoor | MoCap/LeGO-LOAM | ||||
@@
RGB-D/LiDAR-64 | ✓ | Handheld/Car | Indoor/Outdoor |
DSEC [14] | 2 640px 480px | LiDAR-16 | ✗ | Car | Outdoor | RTK-GPS | ||||
AGRI-EBV [15] | 240px 180px | RGB-D/LiDAR-16 | ✓ | Wheeled Robot | Outdoor | LeGO-LOAM | ||||
TUM-VIE [16] | 2 1280px 720px | ✗ | ✓ | Handheld/Bike | Indoor/Outdoor | MoCap | ||||
VECtor [17] | 2 640px 480px | RGB-D/LiDAR-128 | ✓ | Handheld/Scooter | Indoor | MoCap/ICP | ||||
M3ED [18] | 2 1280px 720px | LiDAR-64 | ✓ | Multiple Robots | Indoor/Outdoor | Faster-LIO | ||||
CEAR (ours) |
@@
* Extension -N represents a N-beam LiDAR and -D means depth camera.
@@
Event cameras, inspired by biological vision mechanisms, detect logarithmic intensity changes in images and offer low latency and high temporal resolution [19], making them exceptionally suited for handling rapid robotic movements [20]. Additionally, their high dynamic range (HDR) effectively addresses challenges in poorly-lit and brightly-lit settings, thereby broadening potential applications for robots [21, 22]. Recent studies on event cameras have increasingly gained attention in systems requiring rapid perception, such as drones and autonomous vehicles [23, 24], yet their use has not been widely recognized within the legged robot community. Once properly configured, these unique sensors can significantly enhance state estimation and terrain perception, greatly expanding the capabilities of legged robots.
@@
While promising, event cameras also have several limitations when used as standalone sensing systems. Due to their operating mechanism of detecting changes in brightness, event data inherently depend on the motion of the camera itself (ego-motion), making it challenging to differentiate between the camera’s movement and changes in the external environment. This issue can obscure the clarity of visual data representations, especially when the robot moves slowly or in directions parallel to environmental textures. Overcoming these limitations and enhancing the functionality of agile-legged robots in diverse environments necessitates the development of algorithms for the synergistic integration of multimodal sensors. However, as of 2024, comprehensive datasets are lacking to support this research.
@@
To bridge this gap, we introduce a new dataset collected by a unique sensor suite on an agile quadruped robot, Mini-Cheetah robot [25]. Our system features an event camera, an RGB-D camera, an IMU, a LiDAR, and joint angle sensors. The integration of event and RGB cameras provides the legged robot with enhanced capabilities for both slow and rapid perception. Depth information, crucial for accurate state estimation and navigation, is supplied by the RGB-D camera and LiDAR. The IMU, which can capture high-frequency angular and acceleration data, is inherently immune to disturbances from external dynamic objects. Furthermore, joint angle data can also provide odometry information through forward kinematics.
@@
Our comprehensive dataset includes various distinct robot gaits, including trotting, bounding, and pronking. It also encompasses a wide range of daily indoor and outdoor environments, under varying lighting conditions (e.g., daytime, nighttime, well-lit, dark, under blinking light, and hybrid light setup having bright light sources in dark environments.). This data collection mechanism provides comprehensive scenarios to evaluate the resilience and adaptability of perception systems across various environments.
In summary, the main contributions of this paper are: 1) the inclusion of dynamic motions of a quadruped robot (e.g., trotting, bounding, pronking, and backflipping) to facilitate rapid perception research, 2) comprehensive coverage of the dataset including over 100 sequences gathered in 31 distinct environments under varying lighting conditions and diverse robotic locomotion gaits, and 3) the provision of 6 DoF ground-truth poses from a motion capture system or an advanced SLAM algorithm, along with precise intrinsic, extrinsic, and temporal offset parameters.
@@ II RELATED WORK
In this section, we provide an overview of existing event camera datasets, which are also summarized in Table I. The first event camera dataset, introduced in 2014, featured a eDVS event camera [26] and an RGB-D camera [7]. In 2016, [8] published a dataset using an event camera ( pixels) for visual navigation of a mobile wheeled robot. In the following year, [9] released a comprehensive handheld dataset that includes motion speed information as a reference for investigating the capabilities of an event camera ( pixels). These datasets have played a crucial role in the introduction of event cameras in research communities.
@@
Subsequent developments have led to datasets with higher-resolution event cameras. In 2022, the ViViD dataset [12] was expanded into ViViD++ [13], incorporating a high-resolution event camera ( pixels) for driving scenarios. This update included repeated data collection along the predefined trajectories at different times, categorizing motions as slow, unstable, or aggressive. TUM-VIE dataset [16] includes stereo event camera recordings ( pixels), capturing diverse environments and activities such as walking, running, skating, and biking. Similarly, VECtor dataset [17] features head-mounted and pole-mounted stereo event camera sequences ( pixels) in indoor settings. These enhanced-resolution event camera datasets improved image clarity and details, enabling more precise analysis.
@@
In the evolution of event camera datasets, datasets for specific research purposes have also been developed. The DDD17 [27] and DDD20 [28] datasets, for instance, provide comprehensive metadata including vehicle speed, GPS positioning, and detailed driving dynamics such as steering, throttle, and brake inputs, improving steering prediction accuracy. [29] aimed to tackle the place recognition challenge by mounting a DAVIS346 event camera ( pixels) on a car, capturing data along the same trajectory at different times. [15] created agricultural robotics datasets in varied agricultural settings using a mobile wheeled robot. Additionally, [14] utilized a dual stereo camera setup with a 16-channel Velodyne LiDAR and an RTK-GPS system for automotive driving scenarios. These datasets delve deeper into specific fields and provide the opportunity to explore the potential of event cameras in focused applications.
@@
Another trend of event camera datasets is their application in various robotic platforms. [11] introduced a dataset focused on aggressive drone racing motions. [10] launched the first multi-robot dataset in 2018, featuring indoor and outdoor places with handheld devices, hexacopters, vehicles, and motorcycles. Additionally, [18] unveiled the M3ED dataset, showcasing event cameras across diverse sensor arrays on multiple robotic platforms, including unmanned aerial vehicles, wheeled ground vehicles, and legged robots. Although the M3ED dataset explores event cameras with a quadruped robot, Spot, it did not include the dynamic movements of the robot, such as pronking or bounding gaits. The M3ED dataset features relatively conservative trotting gait only and limited environmental variations. On the other hand, our dataset encompasses various, dynamic quadruped robot gait and motions, multi-modal sensor data including dense depth data, and diverse scenarios gathered in distinct environments under different lighting conditions.
@@ III-A Sensors Overview
The DAVIS346 event camera ( pixels), with a horizontal and vertical field of view and a dB dynamic range, includes a built-in 6-axis IMU and the capability to capture RGB images. The DVXplorer Lite ( pixels), featuring a horizontal and vertical field of view and a dB dynamic range, also includes a built-in 6-axis IMU but cannot output RGB images. Both event cameras feature sub-millisecond latency and operate effectively in lighting conditions up to 100k lux, with 50% of their pixels responding to 80% contrast. The DVXplorer Lite features a dynamic range of approximately 110 dB and functions from 0.3 lux, while the DAVIS346 offers a slightly higher range of 120 dB starting from 0.1 lux. We put an infrared filter in front of both event cameras to block emissions from the RealSense RGB-D camera.
The RealSense camera D455 operates at resolution and , with and fields of view for its RGB and depth cameras, respectively. Timestamps are set at the midpoint of exposure, and auto exposure is enabled to obtain well-exposed images in different lighting conditions. Thus, the timestamp gap between RGB and depth frames varies, typically from to . VectorNav VN-100 is a 9-axis IMU that comprises a 3-axis gyroscope, 3-axis accelerometer, and 3-axis magnetometer at . Additionally, the dataset also includes the orientation data produced by the IMU sensor in quaternion representation. The Velodyne VLP-16 LiDAR delivers highly accurate depth measurements as point clouds at . The LiDAR is mounted atop the Mini Cheetah, ensuring an unobstructed 360 horizontal and vertical FoV of the surrounding environment, with a range of up to 100 meters. The Mini Cheetah’s 12 custom motors feature joint encoders that measure angular position at . For this dataset, data is recorded at .
@@ III-C Time Synchronization
Every time we start the data collection, the Mini-cheetah performs a body pitching swing to produce specific sinusoidal patterns in the sensor data. Since each camera sensor includes a built-in IMU, we compare the angular velocity data from the event and RealSense cameras with the VectorNav IMU, using only the data aligned with the global y-axis. During this process, an event camera is used as a reference and the timestamps of other sensors are offset based on it. Then, the robot’s joint encoder data are synchronized with the event camera by comparing the angular velocity of the knee joint and the built-in IMU of event camera.
@@ IV Dataset
The CEAR dataset contains 106 sequences including 50 indoor, 40 outdoor, and 16 backflip sequences. Our dataset aims to evaluate algorithm performance: 1) across diverse quadruped gaits and lighting conditions via indoor and outdoor sequences, and 2) in agile motions through backflip sequences. The name of each sequence reflects the recording environment, lighting conditions, and the gaits of the quadruped robot. For instance, downtown1_day_trot represents data gathered in a downtown environment during daytime with a trot gait, while classroom_blinking_comb indicates data collected in a classroom environment under blinking light conditions with a combination of trotting, bounding, and pronking gaits. Fig. 4 further illustrates dynamic characteristics inherent to each gait. In addition, we marked the initial positions of the robot’s four feet in each sequence and confirmed its return to these exact positions, providing an additional metric for pose estimation.
@@ IV-A Outdoor Sequences
The outdoor sequences were captured in 10 distinctive outdoor environments, with four sequences per environment. Each set of four sequences includes variations of two different sets of quadruped gaits (trot-only and combined gaits) under different periods of the day (daytime and nighttime), ensuring a comprehensive representation of real-world settings. Daytime sequences feature illumination levels ranging from 40 lux to 80000 lux, while nighttime sequences range from 0.5 lux to 10 lux. Fig. 5 provides an illustration of all outdoor environments and associated challenges such as varied light conditions from overexposure to underexposure, sunny, cloudy, and foggy weather, diverse visual patterns including featureless terrains and distant skies, dynamic objects (e.g., vehicles and pedestrians), and slippery surfaces. We include the estimation result of Faster-LIO algorithm [31] as ground truth because of its good matches between the initial and final poses, as shown in Fig. 6, and superior accuracy on indoor MoCap sequences as evidenced in Table III.